Circular No. 6085

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Postal Address: Central Bureau for Astronomical Telegrams Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A. IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) Phone 617-495-7244/7440/7444 TWX 710-320-6842 ASTROGRAM CAM

MARSDEN@CFA.HARVARD.EDU or GREEN@CFA.HARVARD.EDU (science)

1993 SB A tentative recovery of this object by J. Luu and D. Jewitt with the

University of Hawaii’s 2.2-m reflector on Sept. 10 has been confirmed by Jewitt and J. Chen on Sept. 29 and by M. R. Kidger and G. V. Williams on images obtained with the 2.5-m Isaac Newton reflector on Aug. 15 and 16. Computations by B. G. Marsden show that this is the first of the transneptunian objects found during the past two years actually known to have a Neptune-crossing orbit, a general solution yielding an obviously unstable result with a = 42.9 AU, e = 0.42, q = 25.0 AU, i = 2^{o}. However, the uncertainty is such that the available observations can be represented almost equally well by an orbit with a = 39.4 AU, e = 0.32, q = 26.8 AU, i = 2^{o}, which is in 2:3 resonance with Neptune and prevents the object from approaching within 21 AU of Neptune. Further details are on MPEC 1994-S06. As with 1993 RO and SC (cf. IAUC 6076), further observations, at this and future oppositions, will be required before it can be proven whether there can be close encounters with Neptune or not.

RE J1255+266 M. Dahlem, Johns Hopkins University; and H.-C. Kreysing, Astro-

nomical Institute, University of Tübingen, report the serendipitous detection with the ROSAT WFC of an extremely bright EUV source at α = 12^{h}55^{m}07^{s}.6, δ = +26^{o}41′21′′ (equinox 2000.0, uncertainty ±20′′). The observations started on June 25.51 UT, when the measured countrate was ∼ 14 cts/s. This temporarily made RE J1255+266 the brightest EUV source in the sky. During the observing run, which ended on July 7.19 UT, the lightcurve showed an exponential decrease in flux with a decay time of ∼ 1 day. At the time of the detection the source was > 3000 times brighter than the sensitivity limit of the ROSAT WFC all-sky survey, during which it was not detected. The Palomar Sky Survey plates show no obvious optical counterpart. A few faint objects close to the source position might belong to the cluster Abell 1656.

PERIODIC COMET BORRELLY (1994l) Total visual magnitude estimates: Sept. 11.43 UT, 10.7 (C. S. Morris,

Pine Mountain Club, CA, 0.26-m reflector); 17.15, 10.5 (A. Diepvens, Balen, Belgium, 0.15-m refractor); 24.46, 10.1 (Morris); 27.12, 10.6 (J. Lancashire, Cambridge, England, 0.30-m refractor).

1994 October 1 Brian G. Marsden